OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y =1/5 x - 10. a. y=5x+10 b.y=5x c..y=-5x-10 d.y=-5x

OpenStudy (anonymous):

@kenyababii123 a line which is perpendicular to other has -ve inverse slope so y=-5x+c it passes through 1,-5 so put x=1and y=-5 -5=-5+c c=0 so answer is (D) :)

OpenStudy (anonymous):

thank you